OpenStudy (anonymous):

What is the equation in point-slope form of the line passing through (0, 5) and (–2, 11)? Please help really struggling here

OpenStudy (anonymous):

You have two points. The two points can gives us the slope. m = (11-5)/(-2-0) m = 6/-2 or -3 Now that you have a slope use it with either point. I'll use the second point (-2, 11) y - (11) = -3(x - (-2) y - 11 = -3(x + 2) y - 11 = -3x - 6 y - 11 = -3x - 6 + 11 y = -3x + 5